Circular No. 7308

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

(216) KLEOPATRA F. Marchis, European Southern Observatory (ESO), La Silla, in collab-

oration with D. Hestroffer, A. Cellino, P. Tanga, and V. Zappala, reports on the direct-imaging observation of (216) Kleopatra, and evidence of the bifurcated shape of this main-belt asteroid: “Observations were done on Oct. 25.27–25.31 UT with the adaptive optics system ADONIS on the 3.6m reflector at La Silla in the K_{s} band (2.15 μm). The observations, which were performed under excellent seeing (0′′.5), seem to rule out wrong (spin) pole solutions and show a double source with two distinct lobes of similar size (flux ratio is ∼ 0.8) and a separation estimated to be 0′′.125. Details can

be found at http://sc6.sc.eso.org/~fmarchis/Science/Kleopatra.”

V1333 AQUILAE S. A. Ilovaisky and C. Chevalier, Observatoire de Haute-Provence,

write: “Observations of V1333 Aql, which includes the optical counterpart of the Aquila X-1 x-ray transient, obtained with the 1.2-m Haute-Provence telescope on Nov. 3–9 UT, reveal that the source apparently returned to its quiescent state sometime between Nov. 8.8 and 9.8 UT. The 6-month duration of this latest outburst (turn-on was reported on May 8; cf. IAUC 7161) is one of the longest observed since the 1990 and 1991 events (IAUC 5281 and 5507).”

NOVA IN THE SMALL MAGELLANIC CLOUD 1999 No. 4 D. L. Welch, McMaster University, writes that the MACHO collabora-

tion (cf. IAUC 6312) has discovered an apparent nova in the Small Magellanic Cloud. Located at α = 0^{h}48^{m}51^{s}.7, δ = −72^{o}30′15′′ (equinox 2000.0), the nova peaked in brightness some time between July 27 and Aug. 3 UT. The last point before eruption was taken on July 27.679, when the object was fainter than V = 19.5. The next observation, on Aug. 3.667, indicates that the nova had brightened by at least 6.5 mag and had saturated the CCD in both V and R. The most recent observation, on Nov. 17.471, shows the nova to be at V = 17.0.

COMET 10P/TEMPEL 2 Visual m_{1} estimates by A. Pearce, Nedlands, W. Australia (0.41-m

reflector): Sept. 12.53 UT, 10.6; 29.53, 10.8; Oct. 10.58, 10.9; 26.51, 11.3; Nov. 1.53, 11.5; 9.53, 12.2.

1999 November 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green